If you own 10,000 shares of a company, how many shares would you own after a...

  1. If you own 10,000 shares of a company, how many shares would you own after a 4-for-1 split?

    A.

    2,500

    B.

    20,000

    C.

    30,000

    D.

    40,000

2 points   

QUESTION 50

  1. Use the following information to calculate a firm's enterprise value:

    Market value of equity: $100 million

    Book value of interest-bearing debt: $30 million

    Excess cash: $10 million

    A.

    $120 million

    B.

    $140 million

    C.

    $160 million

    D.

    $180 million
Answer #1

1. no of share = 10000 * 4 = 40000 (D)

2. Market value of equity 100

Add : Debt 30

Less: cash 10

Enterprise value 120 (A)